Annenbaum Corporation uses the weighted average method in its process costing system. This month, the beginning inventory inAnnenbaum Corporation uses the weighted-average method in its process costing system. This month, the beginning inventory in the first processing department consisted of 2,400 units. The costs and percentage completion of these units in beginning inventory were: Cost Percent Complete Materials costs $ 7,700 65% Conversion costs $ 8,800 45% A total of 10,500 units were started and 7,900 units were transferred to the second processing department during the month. The following costs were incurred in the first processing department during the month: Cost Materials costs $ 127,500 Conversion costs $ 209,000 The ending inventory was 50% complete with respect to materials and 35% complete with respect to conversion costs. What are the equivalent units for conversion costs for the month in the first processing department?
Answer #1

Equivalent units of production = Units transferred out + Ending WIP x Completion
Units started and completed = 7900
Ending WIP (Units) = 2400 + 10500 - 7900 = 5000

Equivalent units for conversion = 7900 + 5000 x 35% = 9650 units

Answer is a. 9650